OpenStudy (anonymous):

Donna purchased 3 bags of dog food.The total cost of her purchase including a tax of 3.96 cost 53.46 what is the cost of one bag A) 3.96 B)13.86 C)16.50 D)1914 Plz plz help and show work plz needed urgently

OpenStudy (anonymous):

Thanks you so uch for trying

OpenStudy (anonymous):

Add 53.46 and 3.94. Then divide it by 3.

OpenStudy (anonymous):

Ok

OpenStudy (anonymous):

But the total with tax and everything is 53.46

OpenStudy (anonymous):

x=cost of a bag of dog food We know that the cost of 3 bags + tax is the total cost, 53.46. We also know that the tax is 3.96. 3x + 3.96 = 53.46 We want to isolate x. 3x + 3.96 - 3.96 = 53.46 - 3.96 3x = 49.50 3x/3 = 49.50/3 x = 16.50 The cost is $16.50.

OpenStudy (anonymous):

Omg thanks do much i had that answer but i didn't know the work